### v1

# Motion Capture from Inertial and Vision Sensors

Abstract
Human motion capture is the foundation for many computer vision and graphics tasks. While industrial motion captures systems with complex camera arrays or expensive wearable sensors have been widely adopted in movie and game production, consumer-affordable and easy-to-use solutions for personal applications are still far from mature. To utilize a mixture of a monocular camera and very few inertial measurement units (IMUs) for accurate multi-modal human motion capture in daily life, we contribute MINIONS in this paper, a large-scale Motion capture dataset collected from INertial and visION Sensors. MINIONS has several featured properties: 1) large scale of over five million frames and 400 minutes duration; 2) multi-modality data of IMUs signals and RGB videos labeled with joint positions, joint rotations, SMPL parameters, etc.; 3) a diverse set of 146 fine-grained single and interactive actions with textual descriptions. With the proposed MINIONS, we conduct experiments on multi-modal motion capture and explore the possibilities of consumer-affordable motion capture using a monocular camera and very few IMUs. The experiment results emphasize the unique advantages of inertial and vision sensors, showcasing the promise of consumer-affordable multi-modal motion capture and providing a valuable resource for further research and development.
Keywords:
Multi-modal Motion Capture Dataset and Benchmark
1 Introduction
Human motion capture is the process of recording human movement represented by a sequence of 3D positions and rotations of mesh or joints of the human body [21, 37, 38]. Industrial motion capture systems have been widely applied in movie and game production, sports analysis, medical diagnosis, etc. However, these systems usually consist of tens of synchronized cameras or a group of wearable sensors and specific signal receivers [1, 2, 50]. Despite their high accuracy for human motion capture, individual consumers can hardly afford the high cost and learn professional configurations. Therefore, this paper investigates accurate multi-modal human motion capture with consumer-affordable devices and easy-to-use operations for daily applications like eXtended Reality (XR), mobile video production, live video streaming, etc.
Figure 1: Overview of our MINIONS dataset. It is collected by multiple types of sensors including eight 2K-resolution RGB cameras, Inertial Measurement Units (IMUs), and an RGB-D scanner. With the multi-modal data, we annotate human motion sequences with (d) 2D/3D joints, (e) the SMPL parameters, (f) the texture of each actor from a scanner, and fine-grained action types with textual descriptions.
Table 1: Comparisons of publicly available human motion datasets for motion capture. MINIONS has more types of actions (Act), more frames, and a longer duration (minutes) of videos with single (S) and multiple (M) actors. Moreover, MINIONS provides both full-body IMU data and HD RGB videos annotated with 3D SMPL Mesh of actors, joints (KP), fine-grained action labels (F-Act), and texture of actors (Texture).
Dataset Year Act Frames Duration Actor Modalities
IMU RGB Mesh KP F-Act Texture
Human3.6M [21] 2014 17 3.6M - S
MuCo-3DHP [38] 2018 8 0.5M - M
MPI-INF-3DHP [37] 2017 8 1.2M - S
3DPW [33] 2018 - 0.05M - M
TotalCapture [50] 2017 5 1.9M 50 S
DIP-IMU [20] 2018 64 0.3M 92 S
HUMAN4D [13] 2020 19 0.05M - M
RICH [19] 2022 - 0.5M - M
BEDLAM [9] 2023 - 0.4M 211 M
H3WB [64] 2023 17 0.1M - S
MINIONS (Ours) 2024 146 5.5M 440 M
Mainstream human motion capture solutions can be divided into marker-based and markerless systems. The former usually requires actors to either wear tight clothes with reflective markers to be captured by an array of infrared cameras or bind a group of IMUs on their bodies that can be sensed by signal receivers [1, 2]. They are expensive and inconvenient for actors to make natural movements. Specifically, IMU-based motion capture has an intrinsic limitation of global location drifting, which will limit long-term motions. The latter mainly consists of dozens of calibrated and synchronized cameras to surround the actors, which not only limits the scope of the actors but also needs complex configurations [21]. In addition, human motion capture from monocular RGB videos has also attracted the attention of researchers due to its cheapness and convenience [24, 31]. Some methods [10, 14] exploit the optimization-based pipeline that fits the motion parameters of Skinned Multi-Person Linear (SMPL) [31] models to human bodies in video frames. Other approaches learn deep Convolutional Neural Networks (CNNs) on several benchmarks to regress the SMPL parameters of human bodies [46]. However, they still suffer from depth ambiguity, occlusions, and quick motion instability in real-life scenarios. Therefore, we aim to explore consumer-affordable multi-modal motion capture using a monocular camera and very few IMUs in daily life.
Despite the existing dataset TotalCapture [50] attempting to address these challenges, it suffers from the lack of variety in scenes, subjects, and actions, as well as the issue of the small scale. Furthermore, actors are required to wear tight-fitting attire throughout the data collection of TotalCapture, leading to inevitable distribution differences between daily life and the experimental environment. These weaknesses limit its potential for widespread application. Therefore, we attempt to build a large-scale dataset that covers diverse common actions performed by single or multiple subjects with daily clothes. The datasets should also contain both videos and IMUs records with accurate human motion annotations like 3D position of joints, and SMPL parameters.
To this end, we contribute a large-scale Motion capture dataset from INertial and visION Sensors, named MINIONS, as shown in Figure 1. The MINIONS dataset has several featured properties: 1) Multi-modalities: it is collected from eight 2K cameras and suits with 17 nine-axis IMU sensors. An RGB-D scanner is also adopted to obtain a textured mesh for each actor. 2) Scalibility: it contains over 5.5 million frames and 440 minutes of action sequences captured from different viewpoints. 3) Diversity: it covers 146 categories of fine-grained actions performed by 36 groups of actors (20 actors for single actions and 16 groups of actors for multi-person interactions). 4) Abundance: it provides abundant time-synchronized annotations for each frame, including 2D/3D joints, SMPL parameters [31], fine-grained action class, and texture of actors. We compare our MINIONS with existing motion capture datasets in Table 1.
Based on MINIONS, we propose a baseline analysis on multi-modal human motion capture with both inertial and vision sensors, and validate the feasibility of stable motion capture using a monocular camera and very few IMUs. More specifically, our experiments reveal that monocular motion capture is often limited by jittering due to blurriness or occlusion, while IMUs-based motion capture tends to suffer from global position drifts caused by casual daily actions and loose-fitting clothes. However, a setup with four to six IMUs and a monocular camera can effectively achieve stable and virtually drift-free motion capture. Our baseline analysis offers substantial experimental evidence that paves the way for further research in this field. Furthermore, MINIONS can also be a benchmark dataset for many other tasks such as 2D-to-3D pose estimation [63, 40], fine-grained action recognition [15, 55], etc. The evaluation results of these tasks further exploit the potential of our dataset.
In summary, the contributions of this paper are three-fold:
  • We build MINIONS, a large-scale human motion dataset from both RGB videos and IMUs, with multi-modal data, multiple actors, diverse actions, and rich annotations for the community.
  • We conduct a detailed baseline analysis of multi-modal human motion capture using both inertial and vision sensors, and further explore the possibilities of consumer-affordable motion capture using a monocular camera and very few IMUs.
  • We provide extensive experiments on mainstream tasks, such as 2D-to-3D pose estimation, fine-grained action recognition, etc., opening up the potential of our datasets.
2 Related Work
Motion Capture by Inertial Sensors [56, 22]. IMUs-based human motion capture systems have been widely applied in the industry because they are robust to changes in illumination and occlusion. Industrial solutions usually exploit full-body IMU sensors which must be bound to human bodies. For example, the Perception Neuron system [1] and the Xsens MVN system [42] both use 17 IMUs for human motion capture. Existing datasets for IMU-based motion capture, e.g., DIP-IMU [20], are collected with such systems. However, the expensive cost, e.g., thousands of dollars for the IMU suit, and complex setup procedures before motion capture greatly raise the application threshold. To reduce the cost of devices and setup, researchers study methods using very few IMUs such as six IMUs. Early studies focused on search-based methods [41, 47]. These methods first recorded the parameters of template actions to build a database. During the inference stage, they used a lazy learning strategy [5] to search for similar actions in the database. However, due to the weak correlation between human motions and sensor parameters, their accuracy and stability are far from applications. Recent approaches mainly adopted optimization-based or regression-based paradigms, which used machine learning or deep learning techniques. For example, optimization-based methods [35, 14] optimized all pose parameters in a sequence to find an optimal pose trajectory that is consistent with the acceleration parameters of sensors. Regression-based methods [20, 57], utilized deep neural networks to learn a mapping from sensor measurements to body positions and joint rotations from pair-wised data. These learning-based methods provide a promising solution for sparse IMUs-based human motion capture. However, they still suffer from location drift due to the cumulative error of IMUs over a long time, e.g., several minutes. Therefore, we consider visual information from RGB videos as supplementary cues for long-term stable motion capture.
Motion Capture from Monocular Video [25, 60, 61]. Motion capture, i.e., 3D pose estimation, from monocular videos has been a popular topic in the computer vision community [30]. Existing methods mainly utilize parameterized models like SCAPE [7], SMPL [31], and GHUM [6] as an intermediate representation of human motion. Motion capture from videos is formulated as the estimation of model parameters, i.e., positions of joints, rotations of joints, and shapes of bodies, from RGB frames. Recent approaches can also be divided into optimization-based methods and regression-based ones. Optimization-based methods [10, 17, 45] optimized parameters of the model, e.g., SMPL, to fit the joints or silhouettes of human bodies in video frames. These methods usually relied on the accuracy of the joint detector while inevitably suffering from the trade-off between prediction accuracy and computation cost. Regression-based methods, e.g., HMR [24] and ROMP [46], learned deep neural networks from large-scale data to regress the model parameters given a single RGB image. Popular datasets for training these methods include Human3.6M [21], MPI-INF-3DHP [37], 3DPW [33], etc., as listed in Table 1. Although these methods have achieved excellent accuracy and real-time performance on these benchmarks, they may still generate poor results and temporal jitters due to occlusion, fast actions, and subtle movements in real-world scenarios. Therefore, this paper aims to exploit sparse IMUs to overcome these challenges.
Motion Capture by Combination Schemes. Motion capture from both videos and IMUs has also attracted the attention of researchers. Existing methods [18, 32, 34, 27, 44] effectively improve the motion capture accuracy by eliminating multiplayer ambiguity in videos and minimization of location drift of IMUs. For example, Gilbert et al. [16] fused multi-channel volumetric data from multi-view cameras and IMU signals to estimate 3D joints. They also built a dataset, TotalCapture [50], as listed in Table 1. However, TotalCapture only contained 50-minute videos of five types of actions performed by five subjects, limiting the scalability of the dataset and related methods. During the data collection of TotalCapture, actors are mandated to wear tight-fitting attire, resulting in unavoidable discrepancies between the attire worn in daily life and that of the experimental setting. Therefore, we focus on building a large-scale motion capture dataset of daily life actions containing both RGB videos and IMUs and providing a more practical motion capture paradigm using monocular videos and sparse inertial signals for daily applications.
Figure 2: Overview of hardware setup. (a) Full-body Perception Neuron Studio [1] with 17 IMUs in total. (b) Scenarios larger than 36 square meters, using 4 to 8 synchronized cameras. (c) A professional RGB-D scanner for texture recovery.
3 The MINIONS Dataset
The construction of the MINIONS dataset consists of 3.1 Hardware Setup, 3.2 Calibration, 3.3 Textured Mesh Reconstruction, and 3.4 Human Motion Capture.
3.1 Hardware Setup
We collect raw data in multiple scenes using four to eight synchronized cameras and full-body IMU suits with $17$ sensors, as shown in Figure 2 (a)(b).
RGB Cameras: High-speed industrial camera XiC MC023CG-SY [4] with a resolution of $1920\times 1200$ and 30fps, is used to capture multi-view RGB videos. For synchronization, all cameras are connected with optical cables and triggered by a stabilized signal clock.
IMUs: Perception Neuron Studio [1] is applied to acquire full-body inertial data including acceleration, angular velocity, and magnetic orientation. Note that the synchronization between all cameras and IMUs is dependent on the calibrated timecode devices.
RGB-D Scanner: A professional PUNE Scanner based on Microsoft’s Kinect [3], as shown in Figure 2 (c), is used to obtain the static human surface for texture and SMPL shape parameters recovery. The scanner can provide a scan accuracy of less than 1mm, with realistic texture maps of $1280\times 1024$ resolution.
Figure 3: Overview of dataset construction. (a) Textured mesh reconstruction with an RGB-D scanner; (b) 3D joints triangulation and tracking from multi-view videos; (c) Human pose from full-body IMUs data; and (d) Motion recovery from inertial and visual results.
3.2 Calibration
Calibration of Cameras. Following the general image-based calibration, we obtain coarse camera intrinsic matrix $K$ and external matrix $R|T$ according to Zhang’s solution [62]. To improve the calibration quality, we manually mark the points on the ground to obtain a fine calibration. The average distance between calibration points and re-projected points is 0.63 pixels on a resolution of $1920\times 1200$.
Calibration of IMUs. The raw sensor data are measured in the IMU local coordinate system and should be aligned with the SMPL coordinate system. So, we transform the raw data following the DIP approach [20]. Through T-pose calibration, as shown in Figure 2 (a), we obtain maps $R^{IS}$:$F^{S}\rightarrow{F^{I}}$, $R^{TI}$:$F^{I}\rightarrow{F^{T}}$ and $R^{TB}$:$F^{B}\rightarrow{F^{T}}$ between sensor coordinate system $F^{S}$, inertial coordinate system $F^{I}$, SMPL coordinate system $F^{T}$, and bones coordinate system $F^{B}$.
Calibration between Cameras and IMUs. By aligning the T-pose skeleton in the camera coordinate system $F^{C}$ and the SMPL coordinate system $F^{T}$, we obtain the transformation between them $R^{TC}$:$F^{C}\rightarrow{F^{T}}$.
3.3 Textured Mesh Reconstruction
As shown in Figure 3 (a), we obtain RGB-D images from the scanner and convert them to point clouds to generate human models $M_{t}$ with textures in a canonical pose [39, 58]. To obtain the static SMPL shape parameters $\beta$, we adjust the parameters $(\beta,\theta)$ of SMPL $M_{s}$ to be close to $M_{t}$ and in the interior of $M_{t}$ through the optimization as in [58],
$E(\beta,\theta)=E_{J}+E_{skin}+E_{cloth}+E_{reg},$ (1)
where $\beta$ and $\theta$ denote the shape and pose parameters of SMPL, $E_{J}$ penalizes the error between the projection and the observed 2D joints, $E_{skin}$ keeps the scan points belonging to the skin close to the model $M_{s}$, $E_{cloth}$ prevents scan points belonging to clothes to be inside the model $M_{s}$, $E_{reg}$ is a priori term to make the results more reasonable.
3.4 Human Motion Capture
The human motion capture workflow includes 1) 2D joint detection; 2) 3D joint triangulation; 3) tracking; and 4) motion recovery. It provides high-quality annotations including 2D and 3D joints, SMPL parameters, and person identity (ID). To ensure the quality of the dataset, each step is double-checked by experienced annotators, and the erroneous results of detection and tracking are adjusted manually.
Figure 4: Example frame of motion recovery with inertial and visual data.
2D Joints Detection. We adopt the HRNet-w48 [51] as the 2D joints detector because of its excellent performance on the 2D joints detection benchmark [23, 54]. After that, we post-process the 2D joints through DarkNet [59] to reduce jitters and improve accuracy. The detection result contains 25 joints $P_{2d}$ of body, face, and feet in the same format as OpenPose [11]. We discard the uncertain joints with low confidence scores below $0.7$.
3D Joints Triangulation. With the camera parameters $(K,R|T)$ and multi-view 2D joints $P_{2d}$, as shown in Figure 3 (b), we select top-$K$ views with the smallest reconstruction error to triangulate the initial 3D joints $\widetilde{P_{3d}}$, instead of all redundant camera views. To guarantee the quality of reconstruction, we discard the views with large reconstruction errors higher than $0.05$ meters.
Tracking. Unlike the workflow of single-subject motion datasets [21, 37, 50], our dataset track the 3D joints $\widetilde{P_{3d}}$ to get correct multi-person motion trajectories. To achieve efficient and accurate tracking, we estimate the corresponding 3D boxes from the 3D joints $\widetilde{P_{3d}}$ and then match these boxes using the SORT algorithm [8] with the max range as ten frames. Each actor has a distinguished ID, and the erroneous tracking results (including ID switches and assigning new IDs) are adjusted with manual double-checking.
Figure 5: Fine-grained Actions. MINIONS contains 121 single-player actions and 25 multi-player actions including common person-person and person-object interactive actions in daily life.
Motion Recovery. Most human motion datasets [21, 37] with single subject directly recover motion sequences from pure vision sensor. However, as shown in Figure 4 (a), occlusion between multiple subjects and objects is inevitable, a serious problem for pure visual recovery. Fortunately, despite the global position drift, we can still obtain accurate motions of actors wearing full-body IMUs under occlusion and varied illumination, as shown in Figure 3 (c) and Figure 4 (b). Therefore, for the joints that cannot be captured accurately by previous visual steps, we fill in the missing joints through full-body IMUs data to obtain the final 3D joints $P_{3d}$ and 3D mesh. In detail, for frames $x_{i}$ whose joints cannot be determined due to occlusions or low illumination, we search simultaneously forward and backward for visual frames $x_{i-n}$ and $x_{i+m}$ with high confidence ($n,m\geq 1$). Based on the high-confidence joints of $x_{i-n}$ and $x_{i+m}$, and IMUs data such as acceleration and velocity from $x_{i-n}$ to $x_{i+m}$, we estimate the ground truth joint locations of frame $x_{i}$. According to 3DPW [33], the accuracy of IMUs decreases as $m$ and $n$ increase over time. However, the values of $m$ and $n$ are typically controllable due to the aid of visual information, thus ensuring the accuracy of our estimation. After the fusion of inertial and visual information, we adopt SMPL [31] as the parameterized representation and fit it to the final 3D joints $P_{3d}$, as shown in Figure 3 (d) and Figure 4 (c). Similar to SMPLify [10], we constrain the SMPL parameters as follows:
$E(\beta,\theta,t)=E_{3d}+E_{reg},$ (2)
where $E_{3d}$ penalizes the error between the generated 3D joints and the final 3D joints $P_{3d}$, $E_{reg}$ is a priori term applied to make the results more reasonable. Specifically, the $E_{3d}$ loss is as follows:
$E_{3d}=\sum w_{3d}\cdot\left\|\mathcal{J}M(\theta,\beta,t)-P_{3d}\right\|_{2}^{2},$ (3)
where $\beta$ denotes the static SMPL shape parameters obtained in 3.3, $\theta$, $t$, and $w_{3d}$ denote the SMPL pose parameters, translation of the camera, and confidence scores of detection, $\mathcal{J}$ is a pre-trained linear regression matrix used to generate 3D joints from the SMPL mesh $M(\theta,\beta,t)$.
Figure 6: Qualitative results from single-subject motion capture data collection. Results of 3D mesh and the corresponding re-projected full-body 2D joints from various views.
3.5 Dataset Statistics
The detailed statistics of the MINIONS are listed in Table 1. Moreover, MINIONS is divided into MINIONS-S and MINIONS-M depending on the number of actors. MINIONS-S is collected for single-actor actions, which contain a total of 4.5 million video frames, 315 minutes of video duration, and 121 categories of fine-grained single actions. In contrast, MINIONS-M captures the interactive actions between multiple actors, containing a total of one million video frames, 125 minutes of video duration, and 25 categories of common person-person and person-object interactive actions in daily life.
The statistic on the number of frames for different action sequences reflects that most action sequences are smaller than 300 frames (6 seconds) and the longest one can be up to 1500 frames (50 seconds), which shows the complexity of human actions in life. We show some examples of our dataset in Figure 5. Please refer to the supplementary materials for more information on fine-grained actions.
Figure 7: Qualitative results from multi-subjects motion capture data collection. Results of 3D mesh and the corresponding re-projected full-body 2D joints from various views. The identities of subjects are distinguished by colors.
4 Experiments
In this section, we first show the qualitative results of our dataset. Then we conduct a detailed baseline analysis of multi-modal human motion capture on our MINIONS and further explore the possibilities of consumer-affordable motion capture using a monocular camera and very few IMUs. At last, we evaluate the mainstream methods on various tasks, such as 2D-to-3D pose estimation and fine-grained video action recognition.
4.1 Qualitative Results
We show the qualitative results of our annotated 3D SMPL mesh and the corresponding re-projected full-body 2D joints from different viewpoints in Figure 6 and Figure 7. Benefiting from the multi-view vision sensors and full-body inertial devices, our pipeline can recover the whole-body human motion and global position under various environmental conditions.
4.2 Multi-modal Human Motion Capture
In this subsection, we perform a baseline analysis of multi-modal human motion capture on our MINIONS dataset and explore the potential for consumer-affordable motion capture using a monocular camera and a minimal number of IMUs.
Dataset. We perform all experiments of this subsection on our MINIONS-S dataset. We divide our dataset into a training set, a validation set, and a testing set by actors. The training set contains 12 actors with 2.2 million frames. The validation set has three actors with 0.9 million frames. The testing set has five actors with 1.4 million frames. To evaluate the performance of motion capture, we use the following metrics: the mean global rotation error $\mu_{glo}$ of all body joints (in degree), the MPJPE of all body joints (in mm), and the average jerk of body joints [57] (in $10^{2}m/s^{3}$).
Baseline Model. To explore consumer-affordable multi-modal motion capture, we utilize the multi-model motion capture model SparseNet to explore the supplementary features from inertial and vision sensors in this subsection. SparseNet is designed based on the real-time monocular motion capture model ROMP [46] and the sparse IMUs-based motion capture model Transpose [57]. This incremental baseline model enables us to emphasize the distinct contributions of each modality, without being obscured due to differences in model architecture and training settings. Please refer to the supplementary materials for the details of our baseline model.
Implementation Details. During training, the input frames of videos are resized to $512\times 512$ and the batch size is set as 512. The model is well-trained for 200 epochs using Adam optimizers with a learning rate of 1e-3. All training and test processes run on an NVIDIA GTX 3090 GPU.
Table 2: Comparation between IMUs-based, monocular vision-based, and multi-modal human motion capture. #IMUs: the number and placement of IMUs used in the algorithms. In detail, 4 means {LeftHand, RightHand, Head, Hip}, 6 means {LeftForeArm, RightForeArm, LeftLeg, RightLeg, Head, Hip}, 6 means {LeftHand, RightHand, LeftFoot, RightFoot, Head, Hip}, 8 means {LeftHand, RightHand, LeftForeArm, RightForeArm, LeftFoot, RightFoot, Head, Hip}, 10 means {LeftHand, RightHand, LeftForeArm, RightForeArm, LeftFoot, RightFoot, LeftLeg, RightLeg, Head, Hip}. #Cams: the number of camera views used in the algorithms. $\mu_{glo}$ and $\sigma_{glo}$: the mean and variance global rotation error of all body joints in degrees. MPJPE: the mean Euclidean distance between the predicted 3D joint positions and the corresponding ground truth joint positions. Jitter: the average jerk of body joints.
#IMUs #Cams $\mu_{glo}$$\downarrow$ [deg] $\sigma_{glo}$$\downarrow$ [deg] MPJPE $\downarrow$ [mm] Jitter $\downarrow$ [$10^{2}m/s^{3}$]
IMUs-based 4 0 14.87 12.27 70.13 10.71
6 0 11.74 8.79 57.31 11.17
6 0 13.58 10.89 64.64 11.03
8 0 11.09 8.11 53.25 11.06
10 0 11.06 8.19 52.71 11.31
Vision-based 0 1 10.31 7.10 45.44 51.19
Multi-modal 4 1 9.82 6.93 41.93 2.82
6 1 9.23 6.22 39.87 2.24
6 1 9.41 6.61 41.91 2.82
8 1 8.89 6.05 39.78 2.69
10 1 8.82 5.98 39.24 2.59
Comparison. We train and evaluate the baseline model based on different settings on our MINIONS dataset, including IMUs-based, monocular vision-based, and multi-modal human motion capture. Following previous works, we apply $\mu_{glo}$ (mean global rotation error), $\sigma_{glo}$ (variance of global rotation error), and MPJPE (Euclidean distance error) to measure the mean error between prediction and ground truth in Euclidean space and angular space. Additionally, we use the Jitter to measure the average jerk of body joints. Our experimental results are detailed in Table 2. The results from our experiments show that motion capture using the monocular camera is frequently hampered by instability resulting from blurriness or occlusion and leads to large Jitter. In contrast, motion capture systems that rely on IMUs are prone to larger errors in global rotations and Euclidean distance, attributable to the inherent unreliability of the IMUs. However, an integrated system combining just four to six IMUs with a monocular camera can successfully ensure stable motion capture with minimal error and virtually no jitter. Furthermore, our experimental results show that a motion capture system equipped with a monocular camera and more than eight IMUs is unnecessary. They provide negligible improvements in global rotation error and Euclidean distance error, while leading to increased jitter and equipment cost.
Figure 8: Visualization comparisons among the vision-based, IMUs-based, and multi-modal human motion capture. The vertexes are colored by the distances to the ground truth positions.
Visualization. To facilitate a more intuitive comparison, we provide visualization results of vision-based, IMUs-based, and multi-modal motion capture in Figure 8. The vertexes of the 3D human body are colored by the distances between the prediction and the ground truth positions. Additionally, we present the global pose error and positional translation for a motion capture sequence in Figure 9.
Figure 9: Visualization. (a): Average angular error (in degree) over sequences. (b): Average translation error (in mm) over sequences.
The results reveal that the motion capture based on IMUs tends to exhibit a gradual increase in both global pose error and global positional translation over time. However, the integration of a monocular camera and the IMUs can significantly reduce global pose error and also mitigate the inherent global positional drift caused by IMUs. Please refer to the supplementary materials for more visualization.
4.3 Benchmarks on other Tasks
In this subject, we evaluate mainstream baselines and benchmarks on various research tasks including 2D-to-3D pose estimation [36, 40, 63] and Fine-grained Video Action Recognition [55, 26, 52] , which facilitate more comprehensive explorations of our dataset.
Table 3: 2D-to-3D pose estimation. #Num: the receptive field of input frames.
Method #Num MPJPE $\downarrow$ (mm) PA-MPJPE $\downarrow$ (mm)
Val Test Val Test
Simple3d [36] 1 24.99 26.59 17.40 20.69
Video3d [40] 27 18.36 19.89 12.30 14.63
81 17.86 19.64 11.90 14.37
243 17.18 19.22 11.40 13.95
MotionBERT [63] 243 16.22 18.75 10.87 13.44
2D-to-3D pose estimation. The 2D-to-3D pose estimation task aims to accurately predict the location of 3D joints from the detected 2D human joints, which is a specific type of 3D pose estimation task. The challenge is that depth information is lost in the projection from 3D joints to 2D joints, making the reverse process of deducing 3D joints from 2D observations ill-posed. In Table 3, we evaluate the accuracy of the mainstream method. MPJPE is applied for measuring the average Euclidean distance between the ground truth and predicted joint positions, while PA-MPJPE (Procrustes-Aligned Mean Per Joint Position Error) is a similar metric that first aligns the predicted pose to the ground truth using a rigid transformation before computing the error. To explore the upper bounds of current mainstream methods, we use the groundtruth of 2D joints as input. The experimental results indicate that there is still room for improvement in this task.
Fine-grained Video Action Recognition. Our dataset provides fine-grained action labels and motion videos from various views, which can be validated for applications in fine-grained video action recognition. As shown in Table 4, we test the accuracy of the mainstream video action understanding methods on our dataset. Following these methods, Top-1 and Top-5 are used to evaluate model performance. Top-1 is the model’s accuracy in predicting the most likely action, while Top-5 is the accuracy at which the true action is within the model’s top-five predictions. Compared to the mainstream video understanding datasets, e.g. Kinetics400 [12], NTU RGB+D 60 [43], and NTU RGB+D 120 [29], our dataset is more challenging since existing methods easily achieve Top-1 accuracy of more than 85% on the Kinetics400 dataset.
5 Discussions
5.1 Ethical Issues
The principal ethical consideration addressed in this paper revolves around the issue of privacy. It is our utmost priority to ensure the protection and confidentiality of the data pertaining to the individuals incorporated in our dataset. We recognize the importance of privacy and are committed to implementing robust measures to safeguard it. In addition, our dataset is not intended for indiscriminate distribution. It is specifically designed and will be exclusively disseminated for the purpose of academic and scholarly research. This will be enforced through the implementation of a rigorous licensing agreement, thereby ensuring that the data is used responsibly and in alignment with our ethical commitment to privacy.
Table 4: Fine-grained Video Action Recognition on our MINIONS dataset.
Method Top-1 Accuracy $\uparrow$ Top-5 Accuracy $\uparrow$
Val Test Val Test
TSN [53] 44.94% 47.74% 77.00% 82.87%
I3D [12] 56.82% 59.64% 88.74% 90.41%
R(2+1)D [49] 55.48% 55.59% 85.78% 88.78%
irCSN [48] 58.51% 61.82% 92.77% 92.67%
SlowFast [15] 53.93% 58.64% 83.82% 88.52%
TSM [28] 59.37% 54.31% 88.64% 85.61%
TPN [55] 64.57% 63.97% 92.11% 91.78%
VideoMAE [52] 73.75% 75.46% 96.01% 96.34%
UniFormerV2 [26] 75.88% 77.65% 96.87% 97.29%
5.2 Limitations
Although the experimental results of multi-modal motion capture show that the fusion of vision and inertial devices significantly improves motion capture accuracy, there still is much room for further research and improvement. Furthermore, the current multi-modal motion capture is not equipped to handle the motion capture of multiple actors, which necessitates the consideration of tracking and matching actors with multi-modal inputs. This constitutes a valuable and promising area of study that merits further scholarly exploration.
6 Conclusion
This paper presents a novel paradigm for consumer-affordable multi-modal motion capture using a monocular camera and sparse IMUs, which offers promising opportunities for personal applications. To support research and applications, we have constructed a large-scale motion capture dataset, called MINIONS, using both inertial and vision sensors. The dataset provides a wide range of representations for human motion, including 3D SMPL models, 2D/3D joints, and action labels, among others. With 5 million frames of 146 fine-grained actions, the dataset is highly scalable for various methods and applications. Moreover, we propose a baseline analysis to learn discriminative representations from IMUs and cameras by extracting supplementary features. The results of experiments underscore the distinct benefits offered by the combination of inertial and vision sensors, highlighting the potential of multi-modal motion capture using a monocular camera and sparse IMUs.